Can someone help me with this problem?
I nned to solve this literal equation for the indicated variable:
D = C - s/n (for s) The indicated variable is Depreciation
Thanks,
Sher

Answer by AnlytcPhil(1806)

Can someone help me with this problem? 
I nned to solve this literal equation for the indicated variable: 
D = C - s/n (for s) The indicated variable is Depreciation 

If you mean:

                     s  
            D = C - ---
                     n

Clear of fractions by multiplying every term by LCD = n

           Dn = Cn - s

Add s to both sides

       s + Dn = Cn

Add -Dn to both sides

            s = Cn - Dn

-----------------------------------

If you mean:

                 C - s  
            D = -------
                   n

Clear of fractions by multiplying both sides by LCD = n

           Dn = C - s

Add s to both sides

       s + Dn = C

Add -Dn to both sides

            s = C - Dn

-----------------------------

Edwin